OpenStudy (anonymous):

Add. Simplify by removing factors of 1 28zx + z-7x ----- ----- z^2-49x^2 z+7x

OpenStudy (anonymous):

What? :)

OpenStudy (anonymous):

Add. simplify by removing factors of 1

OpenStudy (anonymous):

step by step so i can make notes to do next one on my own plz

myininaya (myininaya):

\[\frac{28zx}{(z-7x)(z+7x)}+\frac{z-7x}{z+7x}\]

myininaya (myininaya):

i have to go but u need to find a common denominator

OpenStudy (anonymous):

i really need help someone plz

OpenStudy (anonymous):

i really need help someone plz

OpenStudy (anonymous):

Okay, I see the exercise in mathematical notation now, so I know what to do. :) Here is how it goes... 1) 28zx/(z-7x)(z+7x) (explanation: 28zx remains as is for now, but we have to focus on this part (z-7x)(z+7x) so follow up close with what I'm doing here: the first part (z - 7x)...z and -7x multiplies with the first and second of the second parenthesis like so: (z-7x)(z+7x) = z^2 + 7zx - 7zx - 49x^2 and now you keep solving z^2 - 49x^2, and now you have 28zx/z^2-49x^2 + z-7x/z+7x and now you kind of have to apply what you learned about adding fractions...) You multiply them in cross... 28zx * z + 7x and z^2 - 49x^2... and see what you get.

OpenStudy (anonymous):

@tina-ann: Did I help any? :-)

OpenStudy (anonymous):

yes but could you show ur work

OpenStudy (anonymous):

Please forget everything I wrote earlier. :-) I'm not so sure at the moment how I can do it, but I'll post as soon as I get it solved.